OpenStudy (pphalke):

A home has a triangular backyard. The second angle is 6 degrees more than the first angle. The third angle is 14 degrees more than three times the first angle. Find the angles of the triangular backyard. Plzz helpp

OpenStudy (abdullahsadiqi):

What is the total degrees of all triangular ?

OpenStudy (pphalke):

that's what I asked

OpenStudy (abdullahsadiqi):

Ok it is 180 degrees :P

OpenStudy (pphalke):

so how do you go about solving the problem??

OpenStudy (abdullahsadiqi):

first angel = x , 2nd angle = x+6 , 3rd angel = 3x+14 ,, then x+(x+6) +3x+14 = 180 , now fine the value of X ! @pphalke

OpenStudy (dayakar):